OpenStudy (anonymous):

what is the area of a trangle whose sides are 10,13, and 13

OpenStudy (anonymous):

that triangle is inexistent.

OpenStudy (anonymous):

(18(18-10)(18-13)(18-13))^1/2=60

OpenStudy (anonymous):

you guys, that triangle does not exist!

OpenStudy (anonymous):

why

OpenStudy (anonymous):

can you prve if the triangl is not exist

OpenStudy (anonymous):

miscalculation in first answer hyp = 13 one side = 5 other side is = SQRT(13^2-5^2) = 12 ANSWER IS 60

OpenStudy (anonymous):

triangle inequality

OpenStudy (anonymous):

the hypotenuse must always be greater than either of the two legs, unless its equilateral, and in this case it is equal to a leg which cannot happen and 10<13 so therefore it does not exist.

OpenStudy (anonymous):

Triangle exists in that the problem never says that the order of the legs is what you are invisioning just states three legs and they are 10, 13, 13

OpenStudy (anonymous):

please read what i wrote above.

OpenStudy (anonymous):

never states the triangle is a right triangle

OpenStudy (anonymous):

it doesnt have to be right.

OpenStudy (anonymous):

This is not a right triangle, it is isosceles But you do need to find the height, This is where the Pythagorean theorem comes in when you solve for the height you will get that h=12 (notice also that the sides of this triangle are a Pythagorean triple {5,12,13}) Now that we have the height just use the formula for area, which is \[A=(1/2) b*h\] therefore \[A=(1/2)10*12\] A=60

OpenStudy (anonymous):

see drawing attached

OpenStudy (anonymous):

have a nice day onto help someone else

OpenStudy (anonymous):

aww poo. I was wrong. Please resume your normal lives. Had a brain fart.

OpenStudy (anonymous):

elizated yes the triangle exists it is an isosceles triangle with a base of 10 and corresponding sides measuring 13, try drawing a picture

OpenStudy (anonymous):

i know, i realized my mistake. i just find it amusing that 2 people gave me a badge for it! :)

OpenStudy (anonymous):

the triangle exists because 13-13<10<13+13 and 13-10<13<10+13 so if you split the triangle in 2 triangles you will have 169-25=144 so the area will be 10∗12/2=60!!!sorry for my english!!

OpenStudy (anonymous):

don't forget about this Some sufficient conditions for a pair of triangles to be congruent are: SAS Postulate: Two sides in a triangle have the same length as two sides in the other triangle, and the included angles have the same measure. ASA: Two interior angles and the included side in a triangle have the same measure and length, respectively, as those in the other triangle. (The included side for a pair of angles is the side that is common to them.) SSS: Each side of a triangle has the same length as a corresponding side of the other triangle. AAS: Two angles and a corresponding (non-included) side in a triangle have the same measure and length, respectively, as those in the other triangle. Hypotenuse-Leg (HL) Theorem: The hypotenuse and a leg in a right triangle have the same length as those in another right triangle. This is also called RHS (right-angle, hypotenuse, side). Hypotenuse-Angle Theorem: The hypotenuse and an acute angle in one right triangle have the same length and measure, respectively, as those in the other right triangle. This is just a particular case of the AAS theorem.

OpenStudy (anonymous):

Thanks did not know these how they are called in english,Btw if a triangle is isosceles the median from the angle between the same length sides is vertical to the third side so there will be formed two new acute triangles!!!

OpenStudy (anonymous):

Any two sides of a triangle must be greater than a third side for a triangle to a valid triangle.